OpenStudy (anonymous):

which of the following is true? a. tan^-1 (1) = - cot^-1 (1) b. tan^-1 (-1)= cot^-1 (1) c. tan^-1 (-1) = cot^-1 (-1) d. tan^-1 (1) = cot^-1 (-1)

OpenStudy (anonymous):

letter C i got it

OpenStudy (anonymous):

i use wolf alpha here every answer make it false unless letter c..i just need a pro who say im wrong or im correct.

OpenStudy (anonymous):

yeh sorry I was afk...give me a few mins

OpenStudy (anonymous):

no its ok =)

OpenStudy (anonymous):

ok, well tan (pi/4)=1 so tan^-1(1)=(pi/4) and similarly, tan(-pi/4)=-1 so tan^-1(-1)=-(pi/4). Also, cot (pi/4)=1 so cot^-1(1)=(pi/4). Similarly, cot(-pi/4)=-1 so cot^-1(-1)=-pi/4. Checking our answers against the above considerations shows that you are correct, c is true.

OpenStudy (anonymous):

thanks help a lot